3(2x - 3) - 6x = 10
Answer by funmath(2933)

3(2x-3)-6x=10
3(2x)+3(-3)-6x=10
6x-9-6x=10
(6-6)x-9=10
0x-9=10
0-9=10
-9=10
You're not crazy, this problem has no solution. When your variables are eliminated, and the two sides of the equation are not equal, there is no solution.
Happy Calculating!!!